# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=true, bottomline=false, skipabove=skipbelow=

### v1

@@ Abstract
By training to predict the next token in an unlabeled corpus, large language models learn to perform many tasks without any labeled data. However, their next-token-prediction objective arguably limits their performance in scenarios that require planning, such as writing a coherent article. In this paper, we train a module for planning the future writing process via a self-supervised learning objective. By conditioning on generated latent plans, our model extends the successful language model formula to more abstract planning in an unsupervised way. Empirically, we demonstrate that our method improves language modeling performance in general, particularly with respect to the text structure. Because our framework uses a planner module that is unsupervised and external to the language model, new planner modules can be trained at large scale and easily be shared with the community.
@@ 1 Introduction
In the context of sequential decision making, planning is the process of determining a decision policy from a model of the environment (Sutton & Barto, 2018). In this paper, we propose a method for learning to plan future writing from unlabeled data. Concretely, as illustrated in Figure 1, we first transform unlabeled text into a sequence of abstract writing actions which we obtain from a clustered text embedding space. We then train an external planner network which predicts the next writing action from the current context. Finally, we integrate the new learned planning capability into a regular LM by finetuning it on the self-supervised next-token-prediction task in a way that allows it to extract relevant information from predicted writing actions. [^1] We will make code to reproduce our experiments available upon acceptance.
@@ 2.3 Language Modeling
Among the important algorithmic improvements, Retrieval Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020) informs next-token generation by querying a large database of text. A special case of RAG, nearest neighbor language models (Khandelwal et al., 2020), retrieve the closest matches to the past based on similarity in a latent space, transform the distances into a probability distribution, and interpolate it with the LM’s distribution without any additional finetuning. Orthogonal to the above, we propose a method that predicts a sentence into the future in order to inform the generation of the next tokens via finetuning from unlabeled data. In Controllable Text Generation (Zhang et al., 2024a), generation is guided by user-provided attributes such as topic, style, or keywords. Our method can be seen as a form of controlled text generation, where guiding (latent) attributes are predicted by an external model rather than provided by the user. Aiming at separating the global context from the local context, topic-guided LMs (Lau et al., 2017; Wang et al., 2019) condition the generation on topic models inferred from the context. Recent results indicate that this does not improve language modeling because well-trained LMs already possess a good awareness of the current topic (Zheng et al., 2023a). In contrast, our planner predicts future writing actions, which often carry information that is structural rather than topical.
@@ 3.2 Predicting Future Writing Actions with a Planner Module
$\mathbf{p}$ is an absolute position embedding to integrate sequence information (Vaswani et al., 2017). Given the embedding matrix $\mathbf{Z}_{i-1}\in\mathcal{Z}=\left(\mathbb{R}^{d}\right)^{+}$ with $\left(\mathbf{Z}_{i-1}\right)_{j}$ representing vector $\mathbf{z}_{j}$, we predict the following writing action $a_{i}$. Given that the embedding is a set of vectors, we employ a simple Transformer encoder with mean pooling and a linear classifier head:
$\mathbf{Z}^{\prime}_{i-1}=\operatorname{Transformer}\left(\mathbf{Z}_{i-1}\right);\quad\mathbf{o}_{i-1}=\frac{1}{i-1}\sum\limits_{j=1}^{i-1}\left(\mathbf{Z}^{\prime}_{i-1}\right)_{j};\quad p_{\mathcal{A}}=\operatorname{softmax}\left(\mathbf{W}_{o}\mathbf{o}_{i-1}+\mathbf{b}\right).$
Since the semantic relationships between actions are encoded in their embeddings $\mathbf{\overrightarrow{a}}\in\mathcal{R}^{d}$ which we obtained from clustering, we initialize the output matrix $\mathbf{W}_{o}\in\mathbb{R}^{\left|\mathcal{A}\right|\times d}$ with action embeddings.
@@ 3.3 Fine-tuning the Language Model Conditioned on Predicted Writing Actions
While it is feasible to finetune the language model with oracle writing actions, we find that this substantially hurts performance (Section 5.1).
@@ 4.2 Experimental Details
In Section 5.2 we compare our external planner against the LM-internal planner by Wang et al. (2023b). Since this model uses finetuning to adapt the LM itself to predict writing actions, a fair comparison of internal and external planner requires adaptation of our method to this scenario, which we describe in detail in Appendix B. Briefly, the adaptation consists of inserting externally predicted writing actions as tokens, which allows every token that follows to condition on it via attention. Moreover, the adaptation allows finetuning of the original LM parameters.
@@ 4.3 Evaluation
The primary objective of our work is to improve language modeling whose standard metric is perplexity $PPL=\exp(\mathcal{L})$. Additionally, we hypothesize that our planning improves the language modeling performance in particular with respect to text structure. To this end, we measure how well model-generated text corresponds to ground-truth text in terms of writing action sequences. Concretely, given a real text $X=x_{1}\dots x_{n}=t_{1}\dots t_{m}$, we first use the trained models to generate text continuations $\hat{X}=t_{1}\dots t_{i}\hat{t}_{i+1}\dots\hat{t}_{m^{\prime}}$ given the prefix $t_{1}\dots t_{i}$ for some $i$. Then, we use the procedure described in Section 3.1 to obtain the true sequence of writing actions $A=a_{i+1}\dots a_{m}$ and the generated one $\hat{A}=\hat{a}_{i+1}\dots\hat{a}_{m^{\prime}}$. Finally, we compare these sequences via the Levenshtein distance (Levenshtein et al., 1966) (also referred to as edit distance), which measure the number of insertions, deletions, and substitutions necessary to make the sequences equal.
@@
While the end-goal of our work is improving language modeling, the way in which we aim to do that hinges on two components. First, how accurately the planner module can predict oracle writing actions, and second how successfully we can finetune the language model to make use of the information provided by the planner. To measure the former in isolation, we report two metrics: accuracy (i.e., percentage of writing actions predicted correctly), and average rank: The planner predicts a score for each possible action. To calculate the average rank, we order the predicted scores of each action from high to low and observe which rank the oracle action has on average.
@@ 4.4 Baselines
Since our method involves finetuning a language model, improved performance over the language model with no finetuning (henceforth called None) is expected due to adaptation to the finetuning dataset (Wikipedia). However, since the goal of this paper is to identify how well the language model is able to make use of the information in the predicted writing actions (called Predicted), we would like to isolate the benefit introduced from predicted writing actions alone for our main experiments. To this end, we employ the Fixed baseline, which is finetuned on the same dataset, but always receives the same (uninformative) writing action regardless of the context. Finally, we employ the Oracle baseline, which always receives the true next writing action, to serve as an upper bound for planners trained on our extracted writing actions.
Wang et al. (2023b) train an LM-internal planner, to which we compare our external planner. To enable a fair comparison that isolates the effect of internality/externality of the planner, we re-implement their approach in our framework. The implementation details can be found in Appendix B.
@@ 5.1 Main Result
Table 1 shows the results for all combinations of using different baselines at training and inference time, respectively. The main result is that the LM trained with planner-predicted actions (Predicted/Predicted) outperforms the Fixed baseline (Fixed/Fixed) by 1.14 PPL for GPT-2 and 0.35 PPL for OLMo, and by 0.18 and 0.09 in terms of Levenshtein distance respectively. The results also show that finetuning the LM on planner-predicted actions rather than oracle actions is crucial to the success of the model, as it decreases the train/test mismatch. However, there remain differences of 1.63 PPL and 0.47 PPL with respect to Oracle (Oracle/Oracle), indicating that less than half of the potential of integrating the information in the writing action is actuated due to a suboptimal planner. Moreover, in Appendix C we find that, while the oracle action is much better than a random choice, it is not optimal. Hence, further perplexity gains could be made by learning to consistently predict those actions that lead to an even lower perplexity than the oracle actions.
Fixed Oracle Predicted None
Training Test PPL $\downarrow$ Edit $\downarrow$ PPL Edit PPL Edit PPL Edit
GPT-2 Small
Fixed 26.69 5.09 30.39 - - - 32.89 5.15
Oracle 28.95 5.063 23.92 - 26.94 4.80
Predicted 27.89 5.09 25.08 - 25.55 4.90
OLMo 1B
Fixed 11.81 4.77 12.38 - - - 13.08 5.31
Oracle 12.94 4.82 10.99 - 11.99 4.62
Predicted 12.54 4.87 11.42 - 11.46 4.68
Table 1: Model evaluation under different training and conditioning scenarios. The improvement of Predicted/Predicted over Fixed/Fixed indicates the extent to which the model learns to make use of the information in the action (H1), and Oracle/Oracle shows the best possible performance under perfect prediction of writing actions.
@@
Conducting a fair comparison between internal and external planner as described in Appendix B, we find that the external planner performs 1.22 PPL better than the internal planner by Wang et al. (2023b) if we restrict finetuning to parameter-efficient finetuning with LoRA, and .81 PPL better in full finetuning.
@@ 5.3 Ablations
In Tables 5 and 6 in the Appendix, we analyze the ten largest clusters qualitatively. We find that they often correspond well to structures typically found in a Wikipedia article. For example, an article about a person typically contains information about a person’s origin, education, and professional life in a certain order, which a planner would be able to predict in terms of writing actions. Some actions, however, are rather topical than structural. For example, cluster #7 contains mentions of Italian people and places. This indicates that our procedure of generating abstract writing actions (see Section 3.1) can be improved further, e.g. through an encoder whose representations contain more structural rather than topical information.
Model Acc. $\uparrow$ Rank $\downarrow$ PPL $\downarrow$ Edit $\downarrow$
full model 24.7 36.4 25.55 4.90
no sent rep 21.9 38.3 25.59 4.91
no P init 24.3 37.3 25.51 4.89
no AD init 24.7 36.4 27.29 5.21
@@
3 Business 2010: Snecma and GE formed CFM Materials as a 50/50 joint venture.
@@ 6 Discussion
The enhanced LM ability of our method can be attributed to a stronger ability to plan their writing structurally (H3), since conditioning on predicted writing actions causes the model to follow the ground truth text structure, which is indicated by a lower Levenshtein distance in comparison to finetuning without a planner (Section 5.1). Our qualitative analysis of clusters (Section 5.3) suggests that writing actions indeed correspond to common structures in the dataset.
@@
However, our study is limited in some ways. First, while we try to show the generality of our method through experiments with a more powerful, medium-sized OLMo 1B model trained on a general-purpose dataset such as Wikipedia, due to computational constraints we cannot train models of even larger scale and strength. Second, while we improve over a standard architecture with a model better suited for predicting future text, there are several obvious avenues for improvement that remain unexplored, e.g., planning multiple text units into the future and using a search method such as beam search or Monte Carlo Tree Search at inference time to improve planning and find the action that optimizes downstream LM performance. We leave this avenue for future work.
@@ Appendix A Hyperparameters and Experimental Details
We will make our implementation available upon acceptance.
@@
All of our experiments are based on a cleaned Wikipedia dump from March of 2022 that is publicly available via Huggingface [^4] https://huggingface.co/datasets/wikipedia. We apply no further preprocessing to the dataset.
@@
To compute the Levenshtein distance, we use the python-Levenshtein package (https://pypi.org/project/python-Levenshtein/) with default settings.
@@
Appendix B Internal vs external planner setup

### v2

@@ Abstract
By training to predict the next token in an unlabeled corpus, large language models learn to perform many tasks without any labeled data. However, their next-token-prediction objective arguably limits their performance in scenarios that require planning, such as writing a coherent article. In this paper, we train a module for planning the future writing process via a self-supervised learning objective. Given the textual context, this planning module learns to predict future abstract writing actions, which correspond to centroids in a clustered text embedding space. By conditioning on these actions, our model extends the successful language model formula to more abstract planning in an unsupervised way. Empirically, we demonstrate that our method improves language modeling performance in general, particularly with respect to the text structure. Because our framework uses a planner module that is unsupervised and external to the language model, new planner modules can be trained at large scale and easily be shared with the community.
@@ 1 Introduction
In the context of sequential decision making, planning is the process of determining a decision policy from a model of the environment (Sutton & Barto, 2018). In this paper, we propose a method for learning to plan future writing from unlabeled data. Concretely, as illustrated in Figure 1, we first transform unlabeled text into a sequence of abstract writing actions which we obtain from a clustered text embedding space. We then train an external planner network which predicts the next writing action from the current context. Finally, we integrate the new learned planning capability into a regular LM by finetuning it on the self-supervised next-token-prediction task in a way that allows it to extract relevant information from predicted writing actions. [^1] The code is available on Github.
@@ 2.3 Language Modeling
Among the important algorithmic improvements, Retrieval Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020) informs next-token generation by querying a large database of text. A special case of RAG, nearest neighbor language models (Khandelwal et al., 2020), retrieve the closest matches to the past based on similarity in a latent space, transform the distances into a probability distribution, and interpolate it with the LM’s distribution without any additional finetuning. Orthogonal to the above, we propose a method that predicts a sentence into the future in order to inform the generation of the next tokens via finetuning from unlabeled data. In Controllable Text Generation (Zhang et al., 2024a; Mudgal et al., 2024; Guo et al., 2022; Deng & Raffel, 2023; Yang & Klein, 2021), generation is guided by user-provided attributes such as topic, style, or keywords. Our method can be seen as a form of controlled text generation, where guiding (latent) attributes are predicted by an external model rather than provided by the user. Aiming at separating the global context from the local context, topic-guided LMs (Lau et al., 2017; Wang et al., 2019) condition the generation on topic models inferred from the context. Recent results indicate that this does not improve language modeling because well-trained LMs already possess a good awareness of the current topic (Zheng et al., 2023a). In contrast, our planner predicts future writing actions, which often carry information that is structural rather than topical.
@@ 3.2 Predicting Future Writing Actions with a Planner Module
$\mathbf{p}$ is an absolute position embedding to integrate sequence information (Vaswani et al., 2017). Given the embedding matrix $\mathbf{Z}_{i-1}\in\mathcal{Z}=\left(\mathbb{R}^{d}\right)^{+}$ with $\left(\mathbf{Z}_{i-1}\right)_{j}$ representing vector $\mathbf{z}_{j}$, we predict the subsequent writing action $a_{i}$. Given that the embedding is a set of vectors, we employ a simple Transformer encoder with mean pooling and a linear classifier head:
$\mathbf{Z}^{\prime}_{i-1}=\operatorname{Transformer}\left(\mathbf{Z}_{i-1}\right);\quad\mathbf{o}_{i-1}=\frac{1}{i-1}\sum\limits_{j=1}^{i-1}\left(\mathbf{Z}^{\prime}_{i-1}\right)_{j};\quad\mathbf{pr}_{\mathcal{A}}=\operatorname{softmax}\left(\mathbf{W}_{o}\mathbf{o}_{i-1}+\mathbf{b}\right).$
Here, $\mathbf{pr}_{\mathcal{A}}$ is the predicted probability distribution over actions. Since the semantic relationships between actions are encoded in their embeddings $\mathbf{\overrightarrow{a}}\in\mathcal{R}^{d}$ which we obtained from clustering, we initialize the output matrix $\mathbf{W}_{o}\in\mathbb{R}^{\left|\mathcal{A}\right|\times d}$ with action embeddings.
@@ 3.3 Fine-tuning the Language Model Conditioned on Predicted Writing Actions
An alternative approach is to finetune the language model with oracle writing actions, which we also evaluate in Section 5.1. One the one hand we expect that tuning with oracle actions will increase the mismatch between training (oracle actions) and testing (predicted actions), but on the other hand the more reliable oracle actions might induce the LM to rely more strongly on them.
@@ 4.2 Experimental Details
In Section 5.2 we compare our external planner against the LM-internal planner by Wang et al. (2023b). Since this model uses finetuning to adapt the LM itself to predict writing actions, a fair comparison of internal and external planner requires adaptation of our method to this scenario, which we describe in detail in Appendix C. Briefly, the adaptation consists of inserting externally predicted writing actions as tokens, which allows every token that follows to condition on it via attention. Moreover, the adaptation allows finetuning of the original LM parameters.
@@ 4.3 Evaluation
The primary objective of our work is to improve language modeling whose standard metric is perplexity $PPL=\exp(\mathcal{L})$.
Surface-level generation evaluation
Because perplexity does not evaluate actual generations of the model, we also report ROUGE-2 (F1) (Lin, 2004) and MAUVE (Pillutla et al., 2021) scores. ROUGE-2 measures the overlap between generated text and real text in terms of bigrams. Concretely, given a real text $X=x_{1}\dots x_{n}=t_{1}\dots t_{m}$, we first use the trained models to generate text continuations $\hat{X}=t_{1}\dots t_{i}\hat{t}_{i+1}\dots\hat{t}_{m^{\prime}}$ given the prefix $t_{1}\dots t_{i}$ for some $i$. We then compare true and generated continuations at different lengths, we report the average of the different lengths in Section 5.1 and the individual results in Appendix B.
MAUVE is a metric designed to evaluate open-ended text generation. It aims to measure the divergence between the probability distribution of the model and the true data distribution. To do so, it takes a set of true articles and generated articles, and deduces a tractable probability distribution from each by embedding and clustering the articles to form histograms. It then calculates a divergence frontier (Djolonga et al., 2020) between the two distributions, which is analogous to a precision-recall curve applied to generative models. The area under the curve is the MAUVE score. Because MAUVE doesn’t require pairwise comparisons, we generate texts unconditionally, i.e. without context. We generate 1024 tokens, which corresponds roughly to the average article length.
Abstract generation evaluation
Additionally, we hypothesize that our planning improves the language modeling performance in particular with respect to text structure. To this end, we measure how well model-generated texts corresponds to ground-truth texts in terms of writing action sequences. We use the procedure described in Section 3.1 to obtain the sequence of writing actions that corresponds to a real or generated text.
We provide two metrics for this: Levenshtein (Levenshtein et al., 1966) distance and latent perplexity (Deng et al., 2022). The Levenshtein distance (also referred to as edit distance) measures the number of insertions, deletions, and substitutions necessary to make the sequences equal. Similar to ROUGE-2, it compares generated text to ground-truth text in a pairwise manner. Hence, we report results for the same text continuations as for ROUGE-2.
For latent perplexity, we train a simple HMM-based critic to model the distribution of sequences of abstract writing actions of ground-truth text. Latent perplexity results as the perplexity of the sequence of abstract writing actions in the generated text. Latent perplexity evaluates open-ended text generation, and we thus report it on the same unconditionally generated texts as MAUVE.
@@
While the end-goal of our work is improving language modeling, the way in which we aim to do that hinges on two components. First, how accurately the planner module can predict oracle writing actions, and second how successfully we can finetune the language model to make use of the information provided by the planner. To measure the former in isolation, we report two metrics: accuracy (i.e., percentage of writing actions predicted correctly), and average rank. The planner predicts a score for each possible action. To calculate the average rank, we order the predicted scores of each action from high to low and observe which rank the oracle action has on average.
@@ 4.4 Baselines
Since our method involves finetuning a language model via the action adapter (cmp. Figure 1(c)), improved performance over the language model with no finetuning (henceforth called None) is expected due to adaptation to the finetuning dataset (Wikipedia). However, since the goal of this paper is to identify how well the language model is able to make use of the information in the predicted writing actions (called Predicted), we would like to isolate the benefit introduced from predicted writing actions alone for our main experiments. To this end, we employ the Fixed baseline, which is finetuned on the same dataset, but always receives the same (uninformative) writing action regardless of the context. Finally, we employ the Oracle baseline, which always receives the true next writing action, to serve as an upper bound for planners trained on our extracted writing actions. The different baselines correspond to different writing actions being provided to the language model, but each baseline always uses the same type during training and testing. For Predicted, however, it is possible to train using either oracle actions (Predicted-OA) or planner-predicted actions (Predicted-PA).
Wang et al. (2023b) train an LM-internal planner, to which we compare our external planner. To enable a fair comparison that isolates the effect of internality/externality of the planner, we re-implement their approach in our framework. The implementation details can be found in Appendix C.
@@ 5.1 Main Result
Table 1 shows the results. The main result is that the LM trained with planner-predicted actions (Predicted-PA) outperforms the Fixed baseline by 1.14 PPL for GPT-2 and 0.35 PPL for OLMo. It also outperforms Fixed on both surface-level (ROUGE-2, MAUVE) and abstract (Latent PPL, Edit) generation metrics. The results also show a trade-off between finetuning the LM on planner-predicted actions (Predicted-PA) and oracle actions (Predicted-OA). Finetuning on oracle actions damages perplexity, as it introduces a mismatch between training and testing data. Despite this, it is sometimes better in terms of generation metrics. An explanation can be found in the different extent to which the resulting LMs actually follow the plans proposed by their planner, i.e., whether they produce sentences that belong to the cluster corresponding to the action used in generating the sentence. We find that Predicted-OA follows the plan about 40% of the time, compared to only 20% for Predicted-PA. This makes sense, as oracle actions are more informative than predicted actions. Hence, there is a trade-off between reducing the train/test mismatch to achieve low perplexity and following the plan to achieve high generation quality. Finding a balance between these two is a promising direction for future work.
We also observe that there remain differences of 1.63 PPL and 0.47 PPL with respect to Oracle. While it is not possible to train a planner that is exactly equivalent to oracle access, this does suggest that a planner that is better at matching oracle actions can further improve language modeling performance. Moreover, in Appendix D we find that, while the oracle action is much better than a random choice, it is not optimal. Hence, further perplexity gains could be made by learning to consistently predict those actions that lead to an even lower perplexity than the oracle actions.
Base-LM Model PPL $\downarrow$ MAUVE $\uparrow$ Latent PPL $\downarrow$ ROUGE-2 $\uparrow$ Edit $\downarrow$
GPT-2 Small Oracle 23.92 - - - -
Baselines
None 32.89 0.20 438.10 0.013 4.26
Fixed 26.69 0.38 352.8 0.015 3.88
Proposed
Predicted-OA 26.94 0.45 91.6 0.019 3.69
Predicted-PA 25.55 0.44 205.9 0.017 3.78
OLMo 1B Oracle/Oracle 10.99 - - - -
Baselines
None 13.08 0.14 374.22 0.019 4.83
Fixed 11.81 0.45 250.90 0.022 3.42
Proposed
Predicted-OA 11.99 0.43 74.2 0.026 3.39
Predicted-PA 11.46 0.56 178.2 0.025 3.31
Table 1: Model evaluation under different training and conditioning scenarios. Predicted-OA was trained with oracle actions, while Predicted-PA was trained with planner-predicted actions. The improvement of Predicted-OA over Fixed over all metrics indicates the extent to which the model learns to make use of the information in the action (H1), and Oracle shows the best possible performance under perfect prediction of writing actions.
@@
Conducting a fair comparison between internal and external planner as described in Appendix C, we find that the external planner performs 1.22 PPL better than the internal planner by Wang et al. (2023b) if we restrict finetuning to parameter-efficient finetuning with LoRA, and .81 PPL better in full finetuning.
@@ 5.3 Ablations
In Tables 6 and 7 in the Appendix, we analyze the ten largest clusters qualitatively. We find that they often correspond well to structures typically found in a Wikipedia article. For example, an article about a person typically contains information about a person’s origin, education, and professional life in a certain order, which a planner would be able to predict in terms of writing actions. Some actions, however, are rather topical than structural. For example, cluster #7 contains mentions of Italian people and places. This indicates that our procedure of generating abstract writing actions (see Section 3.1) can be improved further, e.g. through an encoder whose representations contain more structural rather than topical information.
Model Acc. $\uparrow$ Rank $\downarrow$ PPL $\downarrow$
full model 24.7 36.4 25.55
no sent rep 21.9 38.3 25.59
no P init 24.3 37.3 25.51
no AD init 24.7 36.4 27.29
@@
3 Business 2010: Snecma and GE formed CFM Materials as a 50/50 joint vefnture.
@@ 6 Discussion
The enhanced LM ability of our method can be attributed to a stronger ability to plan their writing structurally (H3), since conditioning on predicted writing actions causes the model to follow the ground truth text structure, which is indicated by a lower Levenshtein distance and latent perplexity in comparison to finetuning without a planner (Section 5.1). Our qualitative analysis of clusters (Section 5.3) suggests that writing actions indeed correspond to common structures in the dataset.
@@
6.1 Limitations
Our study is limited in some ways. First, while we try to show the generality of our method through experiments with a more powerful, medium-sized OLMo 1B model, due to computational constraints, we cannot train models of even larger scale and strength. Moreover, while we trained on the Wikipedia corpus, which is often considered general-purpose, its diversity is still considerably more limited than the datasets modern LLMs are trained with, which typically subsume a snapshot of the web. Finally, since LMs only achieve nontrivial performance on zero- and few-shot tasks after large amounts of pretraining compute, we do not evaluate the benefit of our method on such downstream tasks. In the future, we want to apply our method to an LLM of medium size (e.g. OLMo 7B) finetuned on a diverse dataset (e.g. Dolma (Soldaini et al., 2024)).
Second, while we improve over a standard architecture with a model better suited for predicting future text, there are several obvious avenues for improvement that remain unexplored. Since we obtain abstract writing actions via clustering in a text embedding space, they do not necessarily encode information that is best suited to help language modeling. We view learning actions tailored to this use case as a promising research direction. Moreover, we want to explore planning multiple text units in the future and using a search method such as beam search or Monte Carlo Tree Search at inference time to improve planning and find the action that optimizes downstream LM performance.
@@ Appendix A Hyperparameters and Experimental Details
All of our experiments are based on a cleaned Wikipedia dump from March of 2022 that is publicly available via Huggingface [^4] https://huggingface.co/datasets/wikipedia. We do not apply any further preprocessing to the dataset.
@@
To compute the Levenshtein distance, we use the python-Levenshtein package (https://pypi.org/project/python-Levenshtein/) with default settings. For ROUGE and MAUVE, we used the python packages ‘rouge‘ and ‘mauve-text‘, respectively. For latent perplexity, we made minor changes to the implementation by Deng et al. (2022), available at https://github.com/florianmai/criticize_text_generation.
@@
Table 4: Hyperparameter Settings
Appendix B Detailed ROUGE and Edit Distance Results
Table 5 shows the Rouge-2 and edit distances between true and generated continuations, when cut off at different generation lengths. Because the edit distance scales linearly with the number of tokens, we normalize results: for 128 tokens we report the actual edit distance, for 256 the edit distance divided by 2, etc.
Base-LM Model 128 256 512 1024 2048
GPT-2 ROUGE-2 $\uparrow$
None 0.0125 0.0134 0.0136 0.013 0.0116
Fixed 0.0148 0.0167 0.0165 0.0154 0.01
Predicted-OA 0.0185 0.0205 0.0211 0.0196 0.02
Predicted-PA 0.0166 0.0178 0.0184 0.0169 0.01
Edit $\downarrow$
None 4.82 4.42 4.12 4.01 3.94
Fixed 4.63 4.09 3.76 3.50 3.39
Predicted-OA 4.39 3.88 3.55 3.35 3.25
Predicted-PA 4.45 3.92 3.67 3.50 3.37
OLMo ROUGE-2 $\uparrow$
None 0.0212 0.021 0.0194 0.0168 0.0141
Fixed 0.0242 0.0247 0.023 0.0203 0.0165
Predicted-OA 0.0275 0.0296 0.0286 0.0255 0.0209
Predicted-PA 0.0277 0.0284 0.0273 0.024 0.0192
Edit $\downarrow$
None 5.01 4.68 4.71 4.84 4.92
Fixed 4.21 3.69 3.38 3.07 2.77
Predicted-OA 4.15 3.64 3.29 3.06 2.84
Predicted-PA 4.2 3.64 3.25 2.89 2.55
Table 5: Details of ROUGE-2 and Edit distance results for different numbers of tokens.
Appendix C Internal vs External Planner Setup